### v2

@@ 6.1 Simulated data experiments
For comparing the quality of the restored images, four evaluation metrics are employed, which are the mean peak signal-to-noise ratio (MPSNR), the mean structural similarity index (MSSIM), the mean feature similarity index (MFSIM), and the erreur relative globale adimensionnelle de synthese (ERGAS). Let $\mathcal{X}^{*}$ denote the restored HSI and $\hat{\mathcal{X}}$ denote the clean HSI. Then ${X^{*}_{::i}}$ and $\hat{X}_{::i}$ denote the $i$-th band of the restored HSI and clean HSI, respectively. The MPSNR value is defined as
$\mathrm{MPSNR}=\frac{1}{I_{3}}\sum^{I_{3}}_{i=1}10\log_{10}\left(\frac{\max^{2}(X^{*}_{::i})}{\mathrm{mse}(X^{*}_{::i},\hat{X}_{::i})}\right),$
@@
$\mathrm{MSSIM}=\frac{1}{I_{3}}\sum^{I_{3}}_{i=1}\mathrm{SSIM}(X^{*}_{::i},\hat{X}_{::i})\quad\mbox{ and }\quad\mathrm{MFSIM}=\frac{1}{I_{3}}\sum^{I_{3}}_{i=1}\mathrm{FSIM}(X^{*}_{::i},\hat{X}_{::i}),$
@@
$\mathrm{ERGAS}=100\sqrt{\frac{1}{I_{3}}\sum^{I_{3}}_{i=1}\frac{\mathrm{mse}(X^{*}_{::i},\hat{X}_{::i})}{\mathrm{mean}(X^{*}_{::i})}}.$
@@
For the proposed MLTL2p method, we set algorithm parameters $\alpha_{\mathcal{S}}=0.1$, and $\alpha_{X}=\alpha_{\mathcal{G}}=0.01$ for Algorithm 2. In the first phase, we set model parameters $w_{j}=10^{2}$, $p=0.1$, and $\delta_{nl}=0$; we set $\gamma=0.8$ for case 1 and 2 and $\gamma=1$ for case 3 and 4; for $\mathbf{R}_{g}$, we set $\delta_{g}=1$, block size $[128,128,128]$ and block rank $[102,102,3]$; for $\mathbf{R}_{l}$, we set $\delta_{l}=1$, block size $[32,32,32]$ and block rank $[26,26,2]$. Algorithm 2 stops at iteration $10$. In the second phase, we set model parameters $w_{j}=10^{2}$, $p=0.1$; we set $\gamma=1.76$ for case 1 and 2 and $\gamma=2.2$ for case 3 and 4; for $\mathbf{R}_{g}$, we set $\delta_{g}=3$, block size $[128,128,128]$ and block rank $[102,102,5]$; for $\mathbf{R}_{l}$, we set $\delta_{l}=3$, block size $[32,32,32]$ and block rank $[26,26,3]$; for $\mathbf{R}_{nl}$, we set $\delta_{nl}=60/\operatorname{median}(\mathcal{W}_{\mathbf{R}})$, block size $[36,128,128]$ and block rank $[32,43,5]$. Algorithm 2 stops if $\|\mathcal{L}^{k+1}-\mathcal{L}^{k}\|_{F}/\|\mathcal{L}^{k+1}\|_{F}\leq 0.005$ and $\|\mathcal{S}^{k+1}-\mathcal{S}^{k}\|_{F}/\|\mathcal{S}^{k+1}\|_{F}\leq 0.005$. For the competing methods, the parameters are manually tuned according to the settings provided in their articles to ensure optimal performance.
Table 1: Average numerical results tested on simulated data
Case
Index
Noisy
MLTL2p
LRTFL0
SNLRSF
LRTD
BM4D
@@
1 MPSNR 19.63 39.35 37.02 39.10 30.97 33.04 31.56
MSSIM 0.276 0.969 0.940 0.965 0.807 0.858 0.856
MFSIM 0.613 0.981 0.968 0.978 0.907 0.918 0.926
ERGAS 0.426 0.041 0.057 0.043 0.114 0.117 0.110
2 MPSNR 18.52 39.03 35.99 34.97 30.64 26.00 31.05
MSSIM 0.245 0.967 0.929 0.915 0.797 0.650 0.840
MFSIM 0.580 0.980 0.962 0.953 0.902 0.826 0.918
ERGAS 0.470 0.043 0.064 0.072 0.120 0.214 0.118
3 MPSNR 14.46 35.51 34.91 26.24 29.32 15.82 30.88
MSSIM 0.234 0.929 0.920 0.704 0.782 0.576 0.829
MFSIM 0.572 0.967 0.957 0.856 0.903 0.764 0.916
ERGAS 0.600 0.074 0.086 0.170 0.133 0.540 0.119
4 MPSNR 16.10 36.49 34.64 35.33 26.47 23.99 30.35
MSSIM 0.176 0.945 0.909 0.925 0.579 0.503 0.819
MFSIM 0.513 0.966 0.949 0.955 0.804 0.759 0.906
ERGAS 0.593 0.061 0.099 0.078 0.208 0.325 0.125
Mtime -- 299.0 265.9 302.2 207.7 104.9 0.6
@@
For visual quality comparison, the HSIs restored by different methods are presented in Fig. 5-8, with a subregion marked by a white box and enlarged in a white box. The MLTL2p method achieves state-of-the-art performance for HSI denoising and destriping, while the competing methods seem to fail to remove noises or restore HSIs with high quality. For example, for removing mixed noise, the BM4D method fails to remove stripes even if the HSI only contains few stripes as shown in Fig. 5(6(g)); the SNLRSF and QRNN3D methods may not remove the stripes and dead lines, when the noise percentage is high, for example, in Fig. 7(5(e))(5(h)). Also, regarding the quality of the restoration, the HSIs restored by the LRTFL0 method may contain staircase artifacts, e.g., in Fig. 5(5(d)); the HSIs restored by the LRTD method contain blur and artifacts, e.g., in Fig. 5(5(f)); the HSIs restored by the QRNN3D method look noisy, e.g., in Fig. 5(5(h)).
@@
Figure 5: Comparison of HSIs (R:112, G:70, B:20) restored by different methods from Chikusei in case 1. The PSNR value for each restored HSI: (a) Noisy image (19.63 dB); (c) MLTL2p (ours) (39.57 dB); (d) LRTFL0 (37.05 dB); (e) SNLRSF (38.70 dB); (f) LRTD (31.82 dB); (g) BM4D (33.79 dB); (h) QRNN3D (32.20 dB).
@@
Figure 6: Comparison of HSIs (R:118, G:78, B:38) restored by different methods from Xiong-An in case 2. The PSNR value for each restored HSI: (a) Noisy image (18.52 dB); (c) MLTL2p (ours) (38.82 dB); (d) LRTFL0 (35.74 dB); (e) SNLRSF (34.76 dB); (f) LRTD (31.90 dB); (g) BM4D (26.16 dB); (h) QRNN3D (31.01 dB).
@@
Figure 7: Comparison of HSIs (R:60, G:27, B:17) restored by different methods from Washington DC Mall in case 3. The PSNR value for each restored HSI: (a) Noisy image (14.66 dB); (c) MLTL2p (ours) (35.28 dB); (d) LRTFL0 (34.62 dB); (e) SNLRSF (26.65 dB); (f) LRTD (28.52 dB); (g) BM4D (16.15 dB); (h) QRNN3D (30.01 dB).
@@
Figure 8: Comparison of HSIs (R:107, G:73, B:21) restored by different methods from Xiong-An in case 4. The PSNR value for each restored HSI: (a) Noisy image (16.02 dB); (c) MLTL2p (ours) (36.66 dB); (d) LRTFL0 (34.95 dB); (e) SNLRSF (35.46 dB); (f) LRTD (25.91 dB); (g) BM4D (23.12 dB); (h) QRNN3D (29.54 dB).
@@ 6.2 Real data experiments
In this subsection, we test the proposed method and the competing methods on two real HSI datasets containing mixed noise. The test images are subimages of size $128\times 128\times 128$ randomly obtained from the HYDICE Urban444 http://www.erdc.usace.army.mil/Media/Fact-Sheets/Fact-Sheet-Article-View/Article/610433/hypercube/ $(307\times 307\times 210)$ and EO-1 Hyperion [^5] http://www.lmars.whu.edu.cn/prof_web/zhanghongyan/resource/noise_EOI.zip $(400\times 200\times 166)$. A selected band of the HSI restored by each method is presented in Fig. 9 and Fig. 10 for HYDICE Urban dataset and EO-1 Hyperion dataset, respectively. It can be observed that the proposed MLTL2p method can remove the stripes while preserving the image details. However, the LRTFL0, LRTD, BM4D and QRNN3D methods are unable to eliminate the stripes when the band is contaminated by heavy mixed noise as shown in Fig. 9; and the SNLRSF, LRTD, BM4D and QRNN3D methods remove not only the noise but also some structural details of the HSI as shown in Fig. 10.

### v3

@@ 6.1 Simulated data experiments
For comparing the quality of the restored images, four evaluation metrics are employed, which are the mean peak signal-to-noise ratio (MPSNR), the mean structural similarity index (MSSIM), the mean feature similarity index (MFSIM), and the erreur relative globale adimensionnelle de synthese (ERGAS). Let $\mathcal{X}^{*}$ denote the restored HSI and $\hat{\mathcal{X}}$ denote the clean HSI. Then ${\boldsymbol{X}^{*}_{::i}}$ and $\hat{\boldsymbol{X}}_{::i}$ denote the $i$-th band of the restored HSI and clean HSI, respectively. The MPSNR value is defined as
$\mathrm{MPSNR}=\frac{1}{I_{3}}\sum^{I_{3}}_{i=1}10\log_{10}\left(\frac{\max^{2}(\boldsymbol{X}^{*}_{::i})}{\mathrm{mse}(\boldsymbol{X}^{*}_{::i},\hat{\boldsymbol{X}}_{::i})}\right),$
@@
$\mathrm{MSSIM}=\frac{1}{I_{3}}\sum^{I_{3}}_{i=1}\mathrm{SSIM}(\boldsymbol{X}^{*}_{::i},\hat{\boldsymbol{X}}_{::i})\quad\mbox{ and }\quad\mathrm{MFSIM}=\frac{1}{I_{3}}\sum^{I_{3}}_{i=1}\mathrm{FSIM}(\boldsymbol{X}^{*}_{::i},\hat{\boldsymbol{X}}_{::i}),$
@@
$\mathrm{ERGAS}=100\sqrt{\frac{1}{I_{3}}\sum^{I_{3}}_{i=1}\frac{\mathrm{mse}(\boldsymbol{X}^{*}_{::i},\hat{\boldsymbol{X}}_{::i})}{\mathrm{mean}(\boldsymbol{X}^{*}_{::i})}}.$
@@
For the proposed MLTL2p method, we set algorithm parameters $\alpha_{\mathcal{S}}=0.1$, and $\alpha_{X}=\alpha_{\mathcal{G}}=0.01$ for Algorithm 2. In the first phase, we set model parameters $w_{j}=10^{-2}$, $p=0.1$, and $\delta_{nl}=0$; we set $\gamma=0.8$ for case 1 and 2 and $\gamma=1$ for case 3 and 4; for $\mathscr{R}_{g}$, we set $\delta_{g}=1$, block size $[128,128,128]$ and block rank $[102,102,3]$; for $\mathscr{R}_{l}$, we set $\delta_{l}=1$, block size $[32,32,32]$ and block rank $[26,26,2]$. Algorithm 2 stops at iteration $10$. In the second phase, we set model parameters $w_{j}=10^{-2}$, $p=0.1$; we set $\gamma=1.76$ for case 1 and 2 and $\gamma=2.2$ for case 3 and 4; for $\mathscr{R}_{g}$, we set $\delta_{g}=3$, block size $[128,128,128]$ and block rank $[102,102,5]$; for $\mathscr{R}_{l}$, we set $\delta_{l}=3$, block size $[32,32,32]$ and block rank $[26,26,3]$; for $\mathscr{R}_{nl}$, we set $\delta_{nl}=60/\operatorname{median}(\mathcal{W}_{\mathscr{R}})$, block size $[36,128,128]$ and block rank $[32,43,5]$. Algorithm 2 stops if $\|\mathcal{L}^{k+1}-\mathcal{L}^{k}\|_{F}/\|\mathcal{L}^{k+1}\|_{F}\leq 0.005$ and $\|\mathcal{S}^{k+1}-\mathcal{S}^{k}\|_{F}/\|\mathcal{S}^{k+1}\|_{F}\leq 0.005$. For the competing methods, the parameters are manually tuned according to the settings provided in their articles to ensure optimal performance.
Table 2: Average numerical results tested on simulated data.
Case Index Noisy MLTL2p LRTFL0 SNLRSF LRTD
FGLR
@@
+BM4D
1 MPSNR 19.63 39.35 37.02 39.10 30.97 33.65 31.56
MSSIM 0.276 0.969 0.940 0.965 0.807 0.900 0.856
MFSIM 0.613 0.981 0.968 0.978 0.907 0.946 0.926
ERGAS 0.426 0.041 0.057 0.043 0.114 0.111 0.110
2 MPSNR 18.52 39.03 35.99 34.97 30.64 32.77 31.05
MSSIM 0.245 0.967 0.929 0.915 0.797 0.887 0.840
MFSIM 0.580 0.980 0.962 0.953 0.902 0.940 0.918
ERGAS 0.470 0.043 0.064 0.072 0.120 0.117 0.118
3 MPSNR 14.46 35.51 34.91 26.24 29.32 32.07 30.88
MSSIM 0.234 0.929 0.920 0.704 0.782 0.874 0.829
MFSIM 0.572 0.967 0.957 0.856 0.903 0.934 0.916
ERGAS 0.600 0.074 0.086 0.170 0.133 0.125 0.119
4 MPSNR 16.10 36.49 34.64 35.33 26.47 29.12 30.35
MSSIM 0.176 0.945 0.909 0.925 0.579 0.793 0.819
MFSIM 0.513 0.966 0.949 0.955 0.804 0.900 0.906
ERGAS 0.593 0.061 0.099 0.078 0.208 0.146 0.125
Mtime -- 299.0 265.9 302.2 207.7 107.5 0.6
@@
For visual quality comparison, the HSIs restored by different methods are presented in Fig. 5-8, with a subregion marked by a white box and enlarged in a white box. The MLTL2p method achieves state-of-the-art performance for HSI denoising and destriping, while the competing methods seem to fail to remove noise or restore HSIs with high quality. For example, for removing mixed noise, the BM4D method fails to remove stripes even if the HSI only contains few stripes as shown in Fig. 5(6(g)); the SNLRSF and QRNN3D methods may not remove the stripes and dead lines, when the noise percentage is high, for example, in Fig. 7(5(e))(5(h)). Also, regarding the quality of the restoration, the HSIs restored by the LRTFL0 method may contain staircase artifacts, e.g., in Fig. 5(5(d)); the HSIs restored by the LRTD method contain blur and artifacts, e.g., in Fig. 5(5(f)); the HSIs restored by the QRNN3D method look noisy, e.g., in Fig. 5(5(h)).
@@
Figure 5: Comparison of HSIs (R:112, G:70, B:20) restored by different methods from Chikusei in case 1. The PSNR value for each restored HSI: (a) Noisy image (19.63 dB); (c) MLTL2p (ours) (39.57 dB); (d) LRTFL0 (37.05 dB); (e) SNLRSF (38.70 dB); (f) LRTD (31.82 dB); (g) FGLR+BM4D (33.97 dB); (h) QRNN3D (32.20 dB).
@@
Figure 6: Comparison of HSIs (R:118, G:78, B:38) restored by different methods from Xiong-An in case 2. The PSNR value for each restored HSI: (a) Noisy image (18.52 dB); (c) MLTL2p (ours) (38.82 dB); (d) LRTFL0 (35.74 dB); (e) SNLRSF (34.76 dB); (f) LRTD (31.90 dB); (g) FGLR+BM4D (34.95 dB); (h) QRNN3D (31.01 dB).
@@
Figure 7: Comparison of HSIs (R:60, G:27, B:17) restored by different methods from Washington DC Mall in case 3. The PSNR value for each restored HSI: (a) Noisy image (14.66 dB); (c) MLTL2p (ours) (35.28 dB); (d) LRTFL0 (34.62 dB); (e) SNLRSF (26.65 dB); (f) LRTD (28.52 dB); (g) FGLR+BM4D (30.69 dB); (h) QRNN3D (30.01 dB).
@@
Figure 8: Comparison of HSIs (R:107, G:73, B:21) restored by different methods from Xiong-An in case 4. The PSNR value for each restored HSI: (a) Noisy image (16.02 dB); (c) MLTL2p (ours) (36.66 dB); (d) LRTFL0 (34.95 dB); (e) SNLRSF (35.46 dB); (f) LRTD (25.91 dB); (g) FGLR+BM4D (30.11 dB); (h) QRNN3D (29.54 dB).
@@ 6.2 Real data experiments
In this subsection, we test the proposed method and the competing methods on two real HSI datasets containing mixed noise. The test images are subimages of size $128\times 128\times 128$ randomly obtained from the HYDICE Urban [^4] http://www.erdc.usace.army.mil/Media/Fact-Sheets/Fact-Sheet-Article-View/Article/610433/hypercube/ $(307\times 307\times 210)$ and EO-1 Hyperion [^5] http://www.lmars.whu.edu.cn/prof_web/zhanghongyan/resource/noise_EOI.zip $(400\times 200\times 166)$. A selected band of the HSI restored by each method is presented in Fig. 9 and Fig. 10 for HYDICE Urban dataset and EO-1 Hyperion dataset, respectively. It can be observed that the proposed MLTL2p method can remove the stripes while preserving the image details. However, the LRTFL0, LRTD, BM4D and QRNN3D methods are unable to eliminate the stripes when the band is contaminated by heavy mixed noise as shown in Fig. 9; and the SNLRSF, LRTD, BM4D and QRNN3D methods remove not only the noise but also some structural details of the HSI as shown in Fig. 10.
@@
6.3 Numerical Convergence Analysis
We conduct numerical convergence analysis on the MLTL2p method in the second phase. In particular, we analyze the convergence of the sequence $\{(\mathcal{S}^{k},[\boldsymbol{X}_{i}^{k}]_{g},[\boldsymbol{X}_{i}^{k}]_{l},[\boldsymbol{X}_{i}^{k}]_{nl},[\mathcal{G}^{k}]_{g},[\mathcal{G}^{k}]_{l},[\mathcal{G}^{k}]_{nl},\mathcal{L}^{k})\}$ generated by the P-BCD algorithm in Algorithm 2 for model (32).
First, we analyze the convergence of $\mathcal{L}^{k}$ and three terms $\mathcal{L}_{g}^{k}$, $\mathcal{L}_{l}^{k}$ and $\mathcal{L}_{nl}^{k}$, associated with $[\boldsymbol{X}_{i}^{k}]_{g},[\boldsymbol{X}_{i}^{k}]_{l},[\boldsymbol{X}_{i}^{k}]_{nl},[\mathcal{G}^{k}]_{g},[\mathcal{G}^{k}]_{l}$, and $[\mathcal{G}^{k}]_{nl}$. The terms $\mathcal{L}_{g}^{k}$, $\mathcal{L}_{l}^{k}$ and $\mathcal{L}_{nl}^{k}$ are defined as follows
$\mathcal{L}_{g}^{k}=$ $\mathcal{W}_{g}^{-1}\mathscr{R}_{g}^{\top}([\mathcal{Y}^{k}]_{g}),$
$\mathcal{L}_{l}^{k}=$ $\mathcal{W}_{l}^{-1}\mathscr{R}_{l}^{\top}([\mathcal{Y}^{k}]_{l}),$
$\mathcal{L}_{nl}^{k}=$ $\mathcal{W}_{nl}^{-1}\mathscr{R}_{nl}^{\top}([\mathcal{Y}^{k}]_{l}),$
where
$[\mathcal{Y}^{k}]_{g}=[\mathcal{G}^{k}]_{g}\times_{1}[\boldsymbol{X}_{1}^{k}]_{g}\times_{2}[\boldsymbol{X}_{2}^{k}]_{g}\times_{3}[\boldsymbol{X}_{3}^{k}]_{g},$
$[\mathcal{Y}^{k}]_{l}=[\mathcal{G}^{k}]_{l}\times_{1}[\boldsymbol{X}_{1}^{k}]_{l}\times_{2}[\boldsymbol{X}_{2}^{k}]_{l}\times_{3}[\boldsymbol{X}_{3}^{k}]_{l},$
$[\mathcal{Y}^{k}]_{nl}=[\mathcal{G}^{k}]_{nl}\times_{1}[\boldsymbol{X}_{1}^{k}]_{nl}\times_{2}[\boldsymbol{X}_{2}^{k}]_{nl}\times_{3}[\boldsymbol{X}_{3}^{k}]_{nl}.$
Note that $\mathcal{L}_{g}^{k}$, $\mathcal{L}_{l}^{k}$ and $\mathcal{L}_{nl}^{k}$ represent the restored images obtained from global, local, nonlocal low-rank regularization, respectively, while $\mathcal{L}^{k}$ represents the overall restored image, according to line 9 in Algorithm 2. To analyze the numerical convergence, we compute the relative errors of $\mathcal{L}^{k}$, $\mathcal{L}_{g}^{k}$, $\mathcal{L}_{l}^{k}$, and $\mathcal{L}_{nl}^{k}$, e.g.,
$\frac{\|\mathcal{L}^{k}-\mathcal{L}^{k-1}\|_{F}}{\|\mathcal{L}^{k}\|_{F}}$
for $\mathcal{L}^{k}$. In Fig. 11(11(a)), we plot the relative errors vs the number of iterations and observe that these restored images exhibit similar convergence behaviors.
Second, we analyze the convergence of $\mathcal{S}^{k}$ using both the relative error and the group sparsity rate. In Fig. 11(11(b)), we plot the relative errors vs the number of iterations on the left axis, and the group sparsity rate, defined as the proportion of nonzero mode-$1$ fibers of $\mathcal{S}^{k}$, on the right axis. We observe that $\mathcal{S}^{k}$ converges stably in terms of both relative error and group sparsity rate.
(a) $\mathcal{L}_{g}$, $\mathcal{L}_{l}$, $\mathcal{L}_{nl}$, and $\mathcal{L}$
(b) $\mathcal{S}$
Figure 11: Numerical convergence analysis.
6.4 Ablation Experiments
We conduct ablation experiments to demonstrate the advantages of our proposed two-phase MLTL2p method, which leverages multi-scale low-rankness, over the methods that utilize only one or two scales of low-rankness. We use Case 2 of Xiong-An for testing. Let G denote global low-rankness, L denote local block low-rankness, and NL denote nonlocal self-similar tensor low-rankness. We compare our two-phase MLTL2p method, also referred to as the two-phase G+L+NL method, with five other methods: single-phase G, L, and G+L methods, as well as two-phase G+NL and L+NL methods. In Fig. 12, we plot the PSNR values of the restored images obtained by each method vs the number of iterations. The result shows that two-phase methods outperform single-phase methods, and the multi-scale method outperforms those using only one or two scales.
Figure 12: Ablation experiments.
6.5 Parameter Analysis
We conduct a parameter analysis on several model parameters. We use Case 3 of Washington DC Mall for testing the MLTL2p method in the second phase. First, we test parameters $\gamma$ and $p$ associated with the variable $\mathcal{S}$. The plot of PSNR values vs parameters $\gamma$ and $p$ is presented in Fig. 13(13(a)). The PSNR value achieves the best performance when $\gamma\in[2.2,3.5]$ and $p\in[0.1,0.2]$. Hence, we set $\gamma=2.2$ and $p=0.1$.
Second, we test parameters $w_{j}$ and $n_{3}$ associated with the variable $[\mathcal{G}]_{g},[\mathcal{G}]_{l}$, and $[\mathcal{G}]_{nl}$. Note that in the second phase of the proposed method we set $w_{j}=10^{-2}$ and $n_{3}=5,3,$ and $5$, respectively, for $[\mathcal{G}]_{g},[\mathcal{G}]_{l}$, and $[\mathcal{G}]_{nl}$. Fig. 13(13(b)) represents the plot of PSNR values vs parameters $w_{j}$ and $n_{3}$ offset, where $n_{3}$ offset is the deviation from the selected $n_{3}$ values. The PSNR values remain high when the $n_{3}$ offset is close to zero and $w_{j}\geq 10^{-2}$.
(a) Parameters $\gamma$ and $p$
(b) Parameters $w_{j}$ and $n_{3}$ offset
Figure 13: Parameter analysis.
@@ 7 Conclusions
Acknowledgments
The authors would like to thank the anonymous reviewers and the associate editor for their valuable comments and suggestions, which led to significant improvements in this article.